Suppose that the mean score of a class of 31 students was 80. The 15 male students in the class had a mean score of 70. What was the mean score for the 16 female students?
Answer by stanbon(75887)

Suppose that the mean score of a class of 31 students was 80. The 15 male students in the class had a mean score of 70. What was the mean score for the 16 female students?
----------
Total points for the class: 80*31 = 2480
---
Total points for the boys: 15*70 = 1050
---
mean score for the girls: (2480-1050)/16 = 89.375
==============
Cheers,
Stan H.
==============